Group Selection Page HELP Controls Group ITDM Vacuum WP78 SPB SQS XO ToBeMigrated2XWiki ToBeMigrated2GitlabPages ToBeMigratedByGroup Migrated2Zulip Archived2PDF not2MigrateOrArchive
SCS SXP Migrated2Zulip Photon Commissioning Team XRO DOC HED Migrated2Zulip WP74 MID PSPO_XO XO_Operation SQS Migrated2Zulip FXE Migrated2Zulip DET Sample Environment SPB Migrated2Zulip
OPERATION-2025 OPERATION-2024 OPERATION-2023 OPERATION-2022 OPERATION-2021 OPERATION-2020 OPERATION-2019 SASE1-2018 SASE1-2017 Experiment data
  OPERATION-2025 logbook, Contact: F.Wolff-Fabris, A.Galler, H.Sinn  Not logged in ELOG logo
Message ID: 70     Entry time: 21 May 2025, 22:41
Author: Konstantin Kharitonov 
Group: SPB 
Subject: Shift summary 

Staff: @**Katerina Doerner**  @**Elizabeth Galtry** @**Raphael de Wijn** @**Mohammad Vakili** @**Chan Kim** @**Konstantin Kharitonov** 

**X-ray delivery**
- 9.3 keV, 1400-1600 uJ pulse energy

**Optical laser delivery**
- No

**Achievements/Observations**
-  Found working nozzle, started injection, observed droplets, performed two parameter scans for HV 
- We could not reach the  stable droplet  frequency around 10 Hz  and were ~  fluctuating 9-17 Hz
- We  prepaired  3 more capillaries  with powder for distance calibration and left them in D12

**Issues**
- Some nozzles were leaking


**Instrument status**
- AGIPD HV off.
- AGIPD gate valve closed and pinned.
- Attenuators set to thickest diamond.
- Vacuum valves closed.
- AGIPD breaks engadged
- ns Lasers off
- HIREX  is handled by FXE, but maybe better to check in the morning

- AGIPD IS REMOVED FROM DAQ  TO MEASURE DROPLET TRAINS DON'T FORGET TO PUT IT BACK

ELOG V3.1.4-7c3fd00